differentiate (x+2)sin3x/ x3+cot2x with respect to x

Grade:11

2 Answers

Vijay Mukati
8 years ago
Dear Student,

In this problem you will need to use the formula for differentiation of two functions in division i.e. d/dx (1/2) = 2(d/dx (1)) – 1 (d/dx(2)) / (22) . Also you need to apply chain rule simultaneously.

Thanks.
siva prasad reddy
8 years ago
) . Also you need to apply chain rule simultaneously.Thanks.or 2n this problem you will need to use the formula for differentiation of two functions in division i.e. d/dx (1/2) = 2(d/dx (1)) – 1 (d/dx(2)) / (2